A small plastic pool has a diamter of 8.5 feet. Find the area of the lawn that will be covered by the pool. Use pie = 3.14.
Answer by nerdybill(7384)

A small plastic pool has a diamter of 8.5 feet. Find the area of the lawn that will be covered by the pool. Use pie = 3.14.
.
area = (pi)(d/2)^2
where
pi is 3.14
d is diameter 8.5 feet
.
area = (3.14)(8.5/2)^2
area = (3.14)(4.25)^2
area = (3.14)(18.0625)
area = 56.72 square feet